OpenStudy (anonymous):

I'm stock aging! solve, 2/p+1 - 1/p-1 = 2p/p^2-1 thanks!

OpenStudy (zzr0ck3r):

stock aging is funny

OpenStudy (zzr0ck3r):

\(\frac{2}{p+1}-\frac{1}{p-1}=\frac{2p}{(p^2-1)}\)?

OpenStudy (anonymous):

yes!

OpenStudy (zzr0ck3r):

use the difference of squares then find a comon denom \(\frac{2}{p+1}-\frac{1}{p-1}=\frac{2p}{(p^2-1)}\iff\frac{2}{p+1}-\frac{1}{p-1}=\frac{2p}{(p-1)(p+1)}\iff\\\frac{2(p-1)}{(p+1)(p-1)}-\frac{1(p+1)}{(p-1)(p+1)}=\frac{2p}{(p^2-1)}\iff\\2(p-1)-(p+1))=2p\iff2p-2-p-1=2p\iff-3=p\)